OpenStudy (anonymous):

Please help! I need to pass this. I don't understand it at all. :( how many solutions does the equation have 3(4x-2)=2(6x-3) A. one solution B. no solution C. infinite solutions

OpenStudy (anonymous):

\[3(4x-2)=2(6x-3) \] lets solve it multiply out, what do you get?

OpenStudy (anonymous):

aka use the distributive property on the left and the right

OpenStudy (anonymous):

if you are not sure how to do it, say so and i will show you

OpenStudy (anonymous):

12? i'm not too sure, sorry!

OpenStudy (anonymous):

you get a \(12\) all right lets go slow

OpenStudy (anonymous):

\[a(b+c)=a\times b+a\times c\] so \[3(4x-2)=3\times 4x+3\times (-2)\] or more simply \[3(4x-2)=12x-6\]

OpenStudy (anonymous):

now we do the same thing on the other side of the equal sign

OpenStudy (anonymous):

ohhh that makes sense.

OpenStudy (anonymous):

what do you get when you do it on the right? we are not done yet